# Scalar tidal response of a rotating BTZ black hole

### v1

@@ Abstract
In this work, we study the tidal response of a rotating BTZ black hole to the scalar tidal perturbation. We show that the real component of the tidal response function isn’t zero, indicating that a rotating BTZ black hole possesses non-zero tidal Love numbers. Additionally, we observe scale-dependent behaviour, known as log-running, in the tidal response function. We also conduct a separate analysis on an extremal rotating BTZ black hole, finding qualitative similarities with its non-extremal counterpart. In addition, we present a procedure to calculate the tidal response function of a charged rotating BTZ black hole as well.
@@ 1 Introduction
Einstein’s general theory of relativity describes gravity as a manifestation of the spacetime curvature [1, 2]. It provides various exciting predictions. One of them is a black hole. Various black hole solutions have been found in $3+1$ spacetime dimensions [3, 4, 1]. Similar black hole solutions have also been explored in lower [5, 6] and higher spacetime dimensions [7] than $3+1$. In this paper, we will study about a black hole in $2+1$ spacetime dimensions.
In 1992, M. Bañados, C. Teitelboim, and J. Zanelli showed that in $2+1$ spacetime dimensions, Einstein-Maxwell equation with negative cosmological constant gives a black hole solution [5, 6] (for review see [8]). These solutions are defined by their mass, angular momentum, and charge. Based on the name of the authors, this black hole is famously known as the BTZ black hole. While BTZ black hole has event and Cauchy horizons (for rotating, charge, and charged rotating cases), it does not have a curvature singularity at the origin [8]. The spacetime metric of a rotating BTZ black hole is given by [9],
@@
In this work, we will study the response of a rotating BTZ black hole to the scalar tidal field. The tidal response of a self-gravitating object can be divided into two parts: conservative (tidal deformation) and dissipative (tidal dissipation) [11]. One can associate dimensionless numbers with these parts, known as tidal Love numbers and tidal dissipation numbers, respectively [11, 12]  [^3] Tidal field can be static (time-independent) or dynamic (time-dependent) in nature, thus, the tidal response function is called static or dynamical tidal response function, respectively (similarly, for tidal Love numbers and tidal dissipation numbers) [12].. Various studies suggest vanishing static tidal Love numbers for black holes in $3+1$ dimensional spacetime [13, 14, 15, 16, 17, 18, 19, 11, 20, 21, 22, 23, 24, 25, 26, 27, 28] . In addition, ref. [29, 30] reported non-zero static scalar tidal Love numbers for a 5-dimensional Myers-Perry black hole. Furthermore, tidal Love numbers have been found to be non-zero for braneworld black holes (presence of an extra spatial dimension) [31]  [^4] Apart from that, tidal Love numbers have also been studied for ultra-compact objects [32, 33], black hole in alternative theory of gravity [34], black hole in asymptotically de Sitter spacetime [35], and area-quantized black hole [36].. Thus, the tidal response of black holes has been studied in $3+1$ and higher dimensions; it would be interesting to study the tidal response of a rotating BTZ black hole, which is in $2+1$ dimensional spacetime, and consequently analyze the nature of its tidal Love numbers. In this work, we will consider a scalar tidal field and study the response of a BTZ black hole to it  [^5] Recently, a paper appeared on the arXiv preprint server [37], which also presented the calculation of the scalar tidal response function of a rotating BTZ black hole. There are differences in the procedure adopted in our work and in ref. [37]. In addition, we have also presented a procedure to calculate the tidal response function of a charged rotating BTZ black hole.. These results might also be helpful to understand the hidden symmetries of general relativity [38, 39].
@@
2 Definition of tidal response function
In this section we will briefly review the definition of the tidal response function of a self-gravitating body (for more details, see  [12, 40, 24]). We will also describe the procedure to calculate the scalar tidal response function of a rotating BTZ black hole.
@@ 2 Definition of tidal response function
where $M/r$ is a point particle contribution to the potential, $\mathcal{E}_{lm}$ are tidal moments, and $F_{lm}$ is defined as the response of the self-gravitating body to the tidal field (or tidal response function; for detailed description, see  [12, 40, 11, 24]). In linear perturbation theory, it can be expressed as $F_{lm}=2k_{lm}+i\omega\tau_{0}\nu_{lm}+\mathcal{O}(\omega^{2})$, where $k_{lm}$ and $\nu_{lm}$ are the tidal Love numbers and tidal dissipation numbers, respectively  [^6] Since we will be working in the $2+1$ spacetime dimension, only the azimuthal index ($m$) will be needed in our analysis. Hence, we will denote the response function as $F_{m}$ instead of $F_{lm}$ (similarly $k_{lm}\to k_{m}$ and $\nu_{lm}\to\nu_{m}$). Also, $l$ will only denote the AdS radius onwards.. $\mathcal{R}$ is related to the radial scale of the body. Therefore, we can define the tidal response function as: the ratio of the coefficients of the decaying part (apart from the point particle contribution) and the growing part of the potential is proportional to the tidal response function.
@@
Thus, our modus operandi for this work will be as follows: First, we shall calculate the form of the scalar perturbation field quantity $(\phi)$ from the source-free Klein-Gordon field equation. Since the Klein-Gordon field equation is a second-order differential equation, we will arrive at two linearly independent solutions. Then we will simplify the general solution by using the purely ingoing boundary condition at the event horizon. Finally, we will find out the behaviour of the field quantity $(\phi)$ in the intermediate region and identify the scalar tidal response function from the above-given definition.
In this work, we are interested in calculating the tidal response function of a BTZ black hole, whose spacetime is not asymptotically flat. However, the above stated procedure to calculate the tidal response function is used for an asymptotically flat spacetime. Thus, a question arise: can we use the above stated procedure for a non-asymptotically flat spacetime? The answer is yes, because we are finally interested in the intermediate region, which is not at asymptotically infinity. Therefore, we can use this approach for a non-asymptotically flat spacetime. One can also use the worldline effective field theory and the scattering amplitude calculation to find the tidal response function of a BTZ black hole [22, 35].
@@ 3 Scalar tidal response function of a non-extremal rotating BTZ black hole
which shifts the origin at the event horizon. The radial source-free Klein-Gordon equation transforms to (for elaborate calculation, see A.1)
@@
To calculate the tidal response function, we need to calculate the large $z$ (or large $r$) limit of the above solution (or the behaviour of the solution in the intermediate region), which will be
@@
The response function depends on the radial coordinate due to the $\ln(z)$ term, which can be termed as the log-running (scale dependent) behaviour of the tidal response function [43, 34]. We can define the tidal response function independent of the radial coordinate as well  [^7] Where we have rewritten the 14 as $R(z)\sim\frac{z^{-a}\Gamma(c)}{\Gamma(a+1)\Gamma(c-a)}\left[1+\tilde{F}_{m}z^{-1}\left\{1+\frac{1}{\tilde{F}_{m}}\ln(z)\right\}\right]~{},$ (16) and defined the tidal response function as the coefficient of $z^{-1}\left\{1+(\ln(z)/\tilde{F}_{m})\right\}$ in the square bracket.:
@@
which is a real quantity. Therefore, it implies that the response function will have non-zero real parts as well, and if we define the real part of the response function as tidal Love numbers of the self-gravitating object. It implies that the scalar tidal Love numbers of a non-extremal rotating BTZ black hole are generically non-zero, and it can be expressed as
@@
Similarly, we can also write the expression of the tidal Love numbers corresponding to the tidal response function with log-running behaviour (15).
Thus, the scalar tidal response function and scalar tidal Love numbers of a non-extremal rotating BTZ black hole are non-vanishing in general, but they vanish in some cases:
  • The tidal response function vanishes for $\omega=0=m$. It implies that the static tidal Love numbers of a non-extremal rotating BTZ black hole vanish in an axi-symmetric scalar tidal field. However, they are non-zero in general.
@@
The multiplication factor for a non-rotating BTZ black hole will be $a(1+a-c)=(\omega^{2}l^{4}-m^{2}l^{2})/4r_{+}^{2}$. Thus, the tidal Love numbers for a non-rotating BTZ black hole are also non-zero. However, they vanish for the same cases listed above for the rotating BTZ black hole.
@@ 4 Scalar tidal response function of an extremal rotating BTZ black hole
In the last section, we have calculated the scalar tidal response function of an arbitrarily rotating BTZ black hole, and we found that it also has a real part. However, the calculation done in the last section is not valid for an extremal black hole $(r_{+}=r_{-})$ due to the transformation (7), which diverges for an extremal rotating BTZ black hole. In this section, we will separately analyze the tidal response function for an extremal rotating BTZ black hole.
Since the transformed coordinate $z$ (7) is not well defined for an extremal rotating BTZ black hole, we re-define the coordinate transformation as
@@
where $C_{1}$ and $C_{2}$ are the constants of integration. Now, we can apply the ingoing boundary condition at the event horizon, which will constrain the constants of integration. Near the event horizon ($r=r_{+}$), i.e., $\tilde{z}\rightarrow 0$, the solution becomes
@@
Based on the decomposition $(\phi=R(r)e^{-i\omega t}e^{im\phi})$, for the solution to be purely ingoing at the event horizon, $C_{2}$ must be zero. Now, the solution becomes
@@
The large $\tilde{z}$ (or large $r$) limit of the above solution (or the behaviour of the solution in the intermediate region) is
@@
The multiplicative factor outside the square bracket in the response function ($B/2$) is a real quantity, and there are real terms inside the square bracket as well. Therefore, the response function will have non-zero real parts, which consequently implies that the extremal rotating BTZ black hole has non-vanishing Love numbers. The Love numbers of an extremal rotating BTZ black hole to a scalar tidal field can be written as
@@
Similarly, we can also express the tidal Love numbers corresponding to the tidal response function with log-running behaviour (28).
Hence, the scalar tidal response function and scalar tidal Love numbers of an extremal rotating BTZ black hole are generally non-zero, but they can be zero for some specific cases:
  • The tidal response function is zero for $\omega=0=m$. Thus, static tidal Love numbers of an extremal rotating BTZ black hole also vanish for an axi-symmetric tidal field. However, they are non-zero, in general.
@@ 5 Scalar tidal response function of a charged rotating BTZ black hole
The spacetime metric of a charged rotating black hole has the same form as given in 1, but the functional form of $f(r)$ is different [5, 44, 45, 46]:
@@
For a charged rotating black hole, due to the logarithmic term in $f(r)$, we can’t write $f(r)$ and $\Omega(r)$ as 2, which is
@@
To proceed further, we will solve the radial Klein-Gordon equation in the near-horizon regime (as done in other studies of tidal response function [11, 24, 27]) and calculate the tidal response function from it. Let us consider a coordinate transformation $\tilde{z}=(r^{2}-r_{+}^{2})/r_{+}^{2}$. In this new coordinate system, the event horizon is located at $\tilde{z}=0$. Now, we can write $f(r)$ as
@@
where we have neglected second and higher order terms of $\tilde{z}$ from $-2q^{2}\ln{\left(r/l\right)}$ in $f(r)$ in the intermediate steps. Here, $\tilde{M}$ and $\tilde{l}$ are given by
@@
Since the approximated expressions of $f(r)$ and $\Omega(r)$ (37) have the same form as 2, we can solve the source-free radial Klein-Gordon equation with the procedure followed in 3, where we can define a new coordinate $\bar{z}=(r^{2}-\tilde{r}_{+}^{2})/(\tilde{r}_{+}^{2}-\tilde{r}_{-}^{2})$, and assume $\tilde{r}_{+}\sim r_{+}$. The solution will be the same as given in 3, with some appropriate changes. These changes are: $r_{\pm}$, $M$, and $l$ will be replaced by $\tilde{r}_{\pm}$, $\tilde{M}$, and $\tilde{l}$, respectively. Thus, the form of the scalar tidal response function for a charged rotating BTZ black hole will also be the same as given in 3, with the above-defined replacements ($(\{r_{\pm},\,M,\,l\}\to\{\tilde{r}_{\pm},\,\tilde{M},\,\tilde{l}\})$.

### v2

@@ Abstract
We study the response of a rotating BTZ black hole to the scalar tidal perturbation. We show that the real component of the tidal response function isn’t zero, indicating that a rotating BTZ black hole possesses non-zero tidal Love numbers. Additionally, we observe scale-dependent behaviour, known as log-running, in the tidal response function. We also conduct a separate analysis on an extremal rotating BTZ black hole, finding qualitative similarities with its non-extremal counterpart. In addition, we present a procedure to calculate the tidal response function of a charged rotating BTZ black hole as well.
@@ 1 Introduction
Einstein’s general theory of relativity describes gravity as a manifestation of the spacetime curvature [1, 2]. It provides various exciting predictions. One of them is a black hole. Various black hole solutions have been found in $3+1$ spacetime dimensions [3, 4, 1, 2]. Similar black hole solutions have also been explored in lower [5, 6] and higher spacetime dimensions [7] than $3+1$. In this paper, we will study about a black hole in $2+1$ spacetime dimensions.
In 1992, M. Bañados, C. Teitelboim, and J. Zanelli showed that in $2+1$ spacetime dimensions, Einstein-Maxwell equation with negative cosmological constant gives a black hole solution [5, 6] (for review see ref. [8]). These solutions are defined by their mass, angular momentum, and charge. Based on the name of the authors, this black hole is famously known as the BTZ black hole. While BTZ black hole has event and Cauchy horizons (for rotating, charge, and charged rotating cases), it does not have a curvature singularity at the origin [8]. The spacetime metric of a rotating BTZ black hole is given by [9],
@@
In this work, we will study the response of a rotating BTZ black hole to the scalar tidal field. The tidal response of a self-gravitating object can be divided into two parts: conservative (tidal deformation) and dissipative (tidal dissipation) [11]. One can associate dimensionless numbers with these parts, known as tidal Love numbers (TLNs) and tidal dissipation numbers, respectively [11, 12]. [^3] In the external observer’s frame, the tidal field can be static (time-independent) or dynamic (time-dependent) in nature, thus, the tidal response function is called static or dynamical tidal response function, respectively (similarly, for tidal Love numbers and tidal dissipation numbers) [12]. Various studies suggest vanishing static TLNs for black holes in $3+1$ dimensional spacetime [13, 14, 15, 16, 17, 18, 19, 11, 20, 21, 22, 23, 24, 25, 26, 27, 28] . In addition, refs. [29, 30] reported non-zero static scalar TLNs for a 5-dimensional Myers-Perry black hole. Furthermore, TLNs have been found to be non-zero for braneworld black holes (presence of an extra spatial dimension) [31]. [^4] Apart from that, TLNs have also been studied for ultra-compact objects [32, 33], black hole in the alternative theory of gravity [34], black hole in Schwarzschild de Sitter spacetime [35], AdS black branes [36], and area-quantized black hole [37]. Thus, the tidal response of black holes has been studied in $3+1$ and higher dimensions; it would be interesting to study the tidal response of a rotating BTZ black hole, which is in $2+1$ dimensional spacetime, and consequently analyze the nature of its TLNs. In this work, we will consider a scalar tidal field and study the response of a BTZ black hole to it. [^5] Recently, a paper appeared on the arXiv preprint server [38], which also presented the calculation of the scalar tidal response function of a rotating BTZ black hole. There are differences in the procedure adopted in our work and in ref. [38]. In addition, we have also presented a procedure to calculate the tidal response function of a charged rotating BTZ black hole. These results might also be helpful to understand the hidden symmetries of general relativity [39, 40].
@@
2 Definition of the tidal response function
In this section, we will briefly review the definition of the tidal response function of a self-gravitating body (for more details, see  [12, 41, 24]). We will also describe the procedure to calculate the scalar tidal response function of a rotating BTZ black hole.
@@ 2 Definition of the tidal response function
where $M/r$ is a point particle contribution to the potential, $\mathcal{E}_{lm}$ are tidal moments, and $F_{lm}$ is defined as the response of the self-gravitating body to the tidal field (or tidal response function; for detailed description, see  [12, 41, 11, 24]). In linear perturbation theory, it can be expressed as $F_{lm}=2k_{lm}+i\omega\tau_{0}\nu_{lm}+\mathcal{O}(\omega^{2})$, where $k_{lm}$ and $\nu_{lm}$ are the tidal Love numbers and tidal dissipation numbers, respectively. [^6] Since we will be working in the $2+1$ spacetime dimension, only the azimuthal index ($m$) will be needed in our analysis. Hence, we will denote the response function as $F_{m}$ instead of $F_{lm}$ (similarly $k_{lm}\to k_{m}$ and $\nu_{lm}\to\nu_{m}$). Also, $l$ will only denote the AdS radius onwards. $\mathcal{R}$ is related to the radial scale of the body and $\tau_{0}$ is the viscosity induced time delay. Therefore, the ratio of the coefficients of the decaying part (apart from the point particle contribution) and the growing part of the potential is proportional to the tidal response function.
@@
Thus, our methodology for this work will be as follows: First, we shall calculate the form of the scalar perturbation field quantity $(\phi)$ from the source-free Klein-Gordon field equation. Since the Klein-Gordon field equation is a second-order differential equation, we will arrive at two linearly independent solutions. Then we will simplify the general solution by using the purely ingoing boundary condition at the event horizon. Finally, we will find out the behaviour of the field quantity $(\phi)$ in the intermediate region [^7] By intermediate region, we mean the region far from the black hole, but well within the tidal field [27]. Thus, the intermediate region is defined as a region of large $r$ ($r\gg r_{+}$) such that $(r-r_{+})~{}\omega\ll 1$, where $r$ is a radial distance from the central object (BTZ black hole in this paper) and $\omega$ is the mode frequency of the tidal field. and identify the scalar tidal response function from the above-given definition.
In this work, we are interested in calculating the tidal response function of a BTZ black hole, whose spacetime is not asymptotically flat. However, the above stated definition to calculate the tidal response function is used for an asymptotically flat spacetime. Thus, a question arises: can we use the above written definition for a non-asymptotically flat spacetime? The answer is yes, because the calculation of the tidal response depends on the growing and decaying behaviour (power law behaviour) of the field quantity, not on the asymptotic of the spacetime [35]. Therefore, we can use this approach for a non-asymptotically flat spacetime. One can also use the worldline effective field theory and the scattering amplitude calculation to find the tidal response function of a BTZ black hole [22, 35].
@@ 3 Scalar tidal response function of a non-extremal rotating BTZ black hole
which shifts the origin at the event horizon. However, the above-defined coordinate transformation is not applicable for the extremal rotating BTZ black hole because it diverges for the extremal rotating BTZ black hole ($r_{+}=r_{-}$). Therefore, the calculation performed in this section is only valid for the non-extremal rotating BTZ black hole. In the next section, we will study the tidal response of an extremal rotating BTZ black hole separately.
The radial source-free Klein-Gordon equation transforms to (for elaborate calculation, see A.1)
@@
To calculate the tidal response function, we need to calculate the large $z$ or large $r$ limit of the above solution (behaviour of the solution in the intermediate region), which will be
@@
The response function depends on the radial coordinate due to the $\ln(z)$ term, which can be termed as the log-running (scale dependent) behaviour of the tidal response function [44, 34]. We can define the tidal response function independent of the radial coordinate as well: [^8] Where we have rewritten the 14 as $R(z)\sim\frac{z^{-a}\Gamma(c)}{\Gamma(a+1)\Gamma(c-a)}\left[1+\tilde{F}_{m}z^{-1}\left\{1+\frac{1}{\tilde{F}_{m}}\ln(z)\right\}\right]~{},$ (16) and defined the tidal response function as the coefficient of $z^{-1}\left\{1+(\ln(z)/\tilde{F}_{m})\right\}$ in the square bracket.
@@
which is a real quantity. Therefore, the tidal response function will have non-zero real parts. Since TLNs are related to the real part of the tidal response function, it implies that the scalar TLNs of a non-extremal rotating BTZ black hole are generically non-zero, and it can be expressed as
@@
Similarly, we can also write the expression of the scalar TLNs corresponding to the tidal response function with log-running behaviour (15).
Thus, the scalar tidal response function and scalar TLNs of a non-extremal rotating BTZ black hole are non-vanishing in general, but they vanish in some cases:
  • The tidal response function vanishes for $\omega=0=m$. It implies that the static TLNs of a non-extremal rotating BTZ black hole vanish in an axi-symmetric scalar tidal field. However, they are non-zero in general.
@@
The multiplication factor for a non-rotating BTZ black hole will be $a(1+a-c)=(\omega^{2}l^{4}-m^{2}l^{2})/4r_{+}^{2}$. Thus, the scalar TLNs of a non-rotating BTZ black hole are also non-zero. However, they vanish for the same cases listed above for the rotating BTZ black hole.
@@ 4 Scalar tidal response function of an extremal rotating BTZ black hole
In the last section, we have calculated the scalar tidal response function of an arbitrarily rotating BTZ black hole, and we found that it has a real part. However, the calculation performed in the last section is not valid for an extremal black hole $(r_{+}=r_{-})$ due to the transformation (7), which diverges for an extremal rotating BTZ black hole. In this section, we will separately analyze the tidal response function for an extremal rotating BTZ black hole.
Since the transformed coordinate $z$ (7) is not well defined for an extremal rotating BTZ black hole, we redefine the coordinate transformation as
@@
where $C_{1}$ and $C_{2}$ are the constants of integration. Now, we can apply the ingoing boundary condition at the event horizon to constrain the constants of integration. Near the event horizon ($r=r_{+}$), i.e., $\tilde{z}\rightarrow 0$, the solution becomes
@@
Based on the decomposition $(\phi=R(r)e^{-i\omega t}e^{im\phi})$, for the solution to be purely ingoing at the event horizon, $C_{2}$ must be zero. Thus, the solution becomes
@@
The large $\tilde{z}$ or large $r$ limit of the above solution (behaviour of the solution in the intermediate region) is
@@
The multiplicative factor outside the square bracket in the response function ($B/2$) is a real quantity, and there are real terms inside the square bracket as well. Therefore, the response function will have non-zero real parts, which consequently implies that the extremal rotating BTZ black hole has non-vanishing scalar TLNs. TLNs of an extremal rotating BTZ black hole to a scalar tidal field can be written as
@@
Similarly, we can also express the TLNs corresponding to the tidal response function with log-running behaviour (28).
Hence, the scalar tidal response function and scalar TLNs of an extremal rotating BTZ black hole are generally non-zero, but they can be zero for some specific cases:
  • The tidal response function is zero for $\omega=0=m$. Thus, static TLNs of an extremal rotating BTZ black hole also vanish for an axi-symmetric tidal field. However, they are non-zero in general.
@@
Having calculated the TLNs for non-extremal and extremal rotating BTZ black holes, we would like to see whether the extremal limit of the TLNs for the non-extremal rotating BTZ black hole gives the TLNs for extremal rotating BTZ black hole or not. Since coordinate transformations involved in the calculations for non-extremal and extremal rotating BTZ black holes are different and the tidal Love numbers show log-running behavior (depends on $z$ logarithmically) in both cases, the extremal limit of the TLNs for non-extremal rotating BTZ black hole will not give the TLNs calculated for extremal rotating BTZ black hole.
@@ 5 Scalar tidal response function of a charged rotating BTZ black hole
The spacetime metric of a charged rotating BTZ black hole has the same form as given in 1, but the functional form of $f(r)$ is different [5, 45, 46, 47]:
@@
For a charged rotating BTZ black hole, due to the logarithmic term in $f(r)$, we can’t write $f(r)$ and $\Omega(r)$ as 2, which is
@@
To proceed further, we will solve the radial Klein-Gordon equation for a slowly varying tidal field ($M\omega\ll 1$), and assume the charge ($q$) being very small, such that $q\sim\mathcal{O}(M\omega)$. Now, we can solve the Klein-Gordon equation in the near-zone region $((r-r_{+})~{}\omega\ll 1)$ and calculate the tidal response function from it. This approach has been followed in various other studies of the tidal response of black holes [11, 24, 27]. Let us consider a coordinate transformation $\tilde{z}=(r^{2}-r_{+}^{2})/r_{+}^{2}$. In this new coordinate system, the event horizon is located at $\tilde{z}=0$. Now, we can write $f(r)$ as (see C for intermediate steps)
@@
where we have neglected terms of order $q^{2}\mathcal{O}(\tilde{z}^{2})$ or $\mathcal{O}(M^{2}\omega^{2}\tilde{z}^{2})$ from $-2q^{2}\ln{\left(r/l\right)}$ in $f(r)$ in the intermediate steps. Here, $\tilde{M}$ and $\tilde{l}$ are given by
@@
Since the approximated expressions of $f(r)$ and $\Omega(r)$ have the same form as 2, we can solve the source-free radial Klein-Gordon equation (33) with the procedure followed in 3. We can define a new coordinate $\bar{z}=(r^{2}-\tilde{r}_{+}^{2})/(\tilde{r}_{+}^{2}-\tilde{r}_{-}^{2})$, which will transform 33 to
$R^{\prime\prime}(\bar{z})+\left[\frac{1}{\bar{z}}+\frac{1}{1+\bar{z}}\right]R^{\prime}(\bar{z})+\frac{1}{\bar{z}^{2}(1+\bar{z})^{2}}\left[\left\{\frac{(\omega-m\tilde{\Omega}_{+})^{2}}{4\tilde{\kappa}_{+}^{2}}-\frac{(\omega-m\tilde{\Omega}_{-})^{2}}{4\tilde{\kappa}_{-}^{2}}\right\}\bar{z}+\frac{(\omega-m\tilde{\Omega}_{+})^{2}}{4\tilde{\kappa}_{+}^{2}}\right]R(\bar{z})=0~{},$ (39)
where $\tilde{\kappa}_{\pm}$ and $\tilde{\Omega}_{\pm}$ are:
$\tilde{\kappa}_{\pm}=\frac{\tilde{r}_{+}^{2}-\tilde{r}_{-}^{2}}{\tilde{l}^{2}\tilde{r}_{\pm}}~{},\qquad\tilde{\Omega}_{\pm}=\frac{\tilde{r}_{\mp}}{\tilde{l}\tilde{r}_{\pm}}~{}.$ (40)
The transformed source-free Klein-Gordon equation is same as 8 with $r_{\pm}$, $M$, and $l$ being replaced by $\tilde{r}_{\pm}$, $\tilde{M}$, and $\tilde{l}$. Thus, we can solve the above-written equation by following the procedure described in 3. Assuming $\tilde{r}_{+}\sim r_{+}$, the solution of the 39 will be the same as given in 3, with some appropriate changes given for the differential equation: $r_{\pm}$, $M$, and $l$ will be replaced by $\tilde{r}_{\pm}$, $\tilde{M}$, and $\tilde{l}$, respectively. Thus, the forms of the scalar tidal response function and the scalar TLNs for a charged rotating BTZ black hole, in the regime of very small charges, will also be the same as given in 3, with the above-defined replacements ($(\{r_{\pm},\,M,\,l\}\to\{\tilde{r}_{\pm},\,\tilde{M},\,\tilde{l}\})$.